### v1

# Semantic Latent Motion for Portrait Video Generation

Abstract
Recent advancements in portrait video generation have been noteworthy. However, existing methods rely heavily on human priors and pre-trained generation models, which may introduce unrealistic motion and lead to inefficient inference. To address these challenges, we propose Semantic Latent Motion (SeMo), a compact and expressive motion representation. Leveraging this representation, our approach achieve both high-quality visual results and efficient inference. SeMo follows an effective three-step framework: Abstraction, Reasoning, and Generation. First, in the Abstraction step, we use a carefully designed Mask Motion Encoder to compress the subject’s motion state into a compact and abstract latent motion (1D token). Second, in the Reasoning step, long-term modeling and efficient reasoning are performed in this latent space to generate motion sequences. Finally, in the Generation step, the motion dynamics serve as conditional information to guide the generation model in synthesizing realistic transitions from reference frames to target frames. Thanks to the compact and descriptive nature of Semantic Latent Motion, our method enables real-time video generation with highly realistic motion. User studies demonstrate that our approach surpasses state-of-the-art models with an 81% win rate in realism. Extensive experiments further highlight its strong compression capability, reconstruction quality, and generative potential. Moreover, its fully self-supervised nature suggests promising applications in broader video generation tasks.
1 Introduction
Portrait video generation has gained significant attention due to its numerous applications in gaming, film, and education. With the development of AI generation technologies, there is an increasing demand for more realistic animation effects and faster generation speeds.
Figure 1: Semantic Latent Motion is a compact and expressive representation of the character’s current state. The motion dynamics in the semantic latent space serve as a powerful guiding condition, steering the generative model from the reference image to the target image. Leveraging the compactness of latent motion, we can model long sequences and perform more efficient resoning.
Figure 2: Overview of SeMo. The proposed framework follows a three-step process (abstraction, reasoning, generation) implemented in two training stages. In the first stage, we train a Motion AutoEncoder for abstraction and generation. During the abstraction step, image features are subjected to random masking and compression to obtain a compact and abstract latent motion. In the generation step, latent motion is used as a condition to guide the diffusion decoder in generating the target image from the reference image. In the second stage, we train a Motion Generator for reasoning, which generates motion sequences from audio.
Existing approaches mainly follow two paths. The first path integrates human prior knowledge of motion representation, such as sequences of 3DMM, pose, landmarks, etc. as expressions of motion [9, 4], combined with a generative renderer to produce the final video frames. The second path utilizes pre-trained video generation models, using their strong video generation capabilities and conditional control of cross-attention to learn the mapping from audio to video frames [6, 4, 41]. Many methods combine both of these paths. However, in the first path, human prior knowledge relies on domain-specific heuristics, often failing to capture detailed motions, which results in human videos that lack naturalness and coherence. Moreover, this framework is not adaptable to multiple scenarios (for example, it cannot be used for general video generation, as it is impractical to extract pose sequences for all objects in the scene). In the second path, the complex spatiotemporal attention modeling and multi-step denoising in diffusion models prevent real-time generation and long-context modeling.
We observe that these motion representations ( 3DMM, pose, landmarks ) all express a high-level semantic:“ the positions and attributes of meaningful parts.” They are all abstractions of human knowledge about facial dynamics and are limited by design biases. For instance, we place landmarks on faces because we believe that changes in the positions of the noise, eyes, and mouth corners are crucial for facial dynamics. Similarly, we use skeletons because we believe that skeletal structures are important for body movements. However, these representations are limited by human abstracting capabilities and cannot capture detailed motion. We ask: Can deep learning models abstract a semantic latent representation that better captures detailed face motions by self-supervised learning? We will introduce our motivation from the following two aspects.
1) Representing video motion in a compact and expressive space. Modeling video motion presents significant challenges. Fortunately, facial videos exhibit more constrained motion ranges and patterns compared to general videos. Previous works have demonstrated that models can synthesize realistic images by extracting detailed information from reference images and semantic information from prior knowledge. Human-crafted priors (e.g., facial landmarks) represent compact and semantic motion representations abstracted through domain knowledge. If humans can achieve this, deep learning models could potentially learn superior representations through self-supervision – capturing more compact and detail-preserving motion characteristics. This motivates our exploration of learning semantic latent motion representations that surpass human-designed priors in expressiveness.
Figure 3: Given an audio clip and a reference image, SeMo can generate lifelike video. From top to bottom, we highlight several key features of our model: a. The generated characters exhibit natural head movements and blinking while speaking. b. They display rich motion dynamics with substantial amplitude. c. Model can adapt to different poses. d. Model generalize well across various datasets. e. The generated videos feature remarkably natural lighting and shadow variations.
2) Employing self-supervised semantic representations to guide generation. Current generative models [5, 17, 1, 2, 43, 23, 46, 13] predominantly rely on human-abstracted conditions like text prompts or landmarks. While effective, these representations are inherently limited by human abstraction capabilities and domain-specific biases, often failing to capture subtle motion details. Recent works like RCG [18] and StoryDiffusion[52] have demonstrated that self-supervised latent features can serve as superior generation conditions compared to manual priors. Building on this insight, we propose using learned semantic latent motion as conditional signals for video generation, enabling more precise and detailed motion control.
Motivated by these insights, we design a universal motion representation framework coupled with a simple yet effective video generation pipeline. Our approach consists of two key components: First, we train a Motion Autoencoder, where an encoder compresses each facial video frame into a compact latent motion representation (1D token) using a random masking strategy. The diffusion-based decoder then reconstructs the target frames using reference images as inputs and semantic latent motion as conditional signals. Once the Motion Autoencoder is trained, we extract latent motion representations for downstream tasks. For portrait video generation, we train a diffusion-based generator that synthesizes latent motion sequences directly from audio inputs. The compact latent space enables efficient long-term temporal modeling and preserves comprehensive event information. During inference, we first generate motion sequences from audio and reference images, then synthesize video frames using the Motion Autoencoder’s decoder.
The rigorous reconstruction task encourages our latent motion representation to develop strong descriptive power, effectively capturing essential information within an extremely compact space. Leveraging this representation, our method achieves highly natural and vivid animations, outperforming state-of-the-art methods with an FID score of 26.8 (compared to the 27.4 baseline) and achieving an 81% win rate in realism in a user study.
Beyond these results, we further observe that even when 90% of the video frame content (input to the motion encoder) is masked, our Motion AutoEncoder can still generate coherent and smooth videos. This reveals two critical insights: 1) High-level semantic information sufficiently guides video generation, 2) Video motion modeling requires less information proportion than image generation (contrasting 75% masking ratio in MAE [14]). Moreover, as a general framework, our Motion AutoEncoder demonstrates strong adaptability to generic videos containing limited subjects and simple motions, such as object rotations.
2 Related Works
Motion-Based Video Codec.
Motion-driven video synthesis techniques typically generate target frames by combining reference frames with motion representations. FOMM [30] established a baseline through sparse keypoint detection and affine transformation estimation to derive dense motion fields. LIA [39] advanced this paradigm by decomposing latent vectors and employing optical flow-based warping operations. MRAA [31] extended the framework by separately predicting foreground region motions and background displacements, subsequently rendering through pixel-wise flow manipulation. IMF [10] introduced a fully self-supervised approach that encodes reference and target frames into compact latent tokens. This method progressively reconstructs target frame pixels through a coarse-to-fine process that integrates reference image features at multiple granularity levels. These developments collectively demonstrate the evolution from explicit motion parameterization towards more implicit, learned representations that better preserve spatial-temporal coherence.
Portrait Video Generation.
The portrait generation task has garnered significant attention in recent years due to its broad applications and has witnessed rapid advancements. Wave2Lip [25] pioneered the use of GAN networks combined with lower-face masking to directly synthesize video frames from reference images and audio inputs. Subsequent works including VividTalk [33], SadTalker [48], and Moda [22] have leveraged 3D morphable priors coupled with neural rendering techniques to generate target video frames. Notably, VASA-1 [42] and Gaia [15] introduced hybrid architectures that encode both 3D prior knowledge and appearance features into a latent space, establishing a more universal and compact motion representation framework. Next, diffusion-based generators are employed to produce motion sequences conditioned on audio inputs [34, 19, 32, 49]. Recent innovations like EMO [34] and Hallo2 [6] have further enhanced visual quality through the integration of pre-trained video generation foundation models with specialized reference networks. AniTalker [21] presents a two-stage self-supervised learning paradigm. The overarching trend in this domain demonstrates a clear shift from hand-crafted priors towards data-driven approaches that enable models to capture finer details and enhance the naturalness of synthesized visuals.
Self-supervised Representation as Guidance.
A large amount of self-supervised learning work  [14, 28, 35, 11, 26, 44, 37] has utilized encoded features for various downstream visual understanding tasks. Self-supervised representations have been proven to possess rich semantics and can serve as powerful guidance signals. In RCG  [18], the quality of image generation was significantly improved by introducing features obtained through MoCo’s [3] self-supervised learning framework. Meanwhile, Story Diffusion [52] obtains a continuous guidance signal by performing linear interpolation in the latent space of CLIP features, thereby controlling the video generation model to produce more coherent content.
3 Method
Figure 4: Illustration of Motion AutoEncoder
Figure 5: An example of video reconstruction: In the example above, we demonstrate that SeMo can not only accurately model facial movements but also capture environmental changes. With a token size of 1, semantic latent motion is not only more compact than the motion in other works but also captures more comprehensive information.
Given a reference image and its corresponding audio, our goal is to generate a dynamic video with synchronized motion and sound. To achieve this, we propose a two-stage self-supervised learning framework. In the first stage, the objective is to extract a compact, high-level semantic motion $\mathbf{M}\in\mathbb{R}^{l\times d}$ for each frame of the image. We first use learnable tokens and random masking to compress the motion into an abstract representation, and then utilize the reference image and the motion to reconstruct the target image. Once the Motion AutoEncoder is trained, it is frozen. In the second stage, we train a Motion Generator. For each frame, we first use the motion encoder to extract the motion and then generate the target motion based on the audio and the corresponding motion of the reference image. The details of these two stages will be described below.
3.1 Motion AutoEncoder
Motion Encoder. Our motion encoder design is motivated by the observation that facial movements - akin to 3D Morphable Models (3DMM) and keypoint trajectories - can be effectively represented in a highly compact space. We propose a novel Masked Motion Encoder architecture that integrates dual strategies of random masking and latent space compression to derive semantically rich latent motion.
Specifically, we first compress input frames $\mathbf{I}\in\mathbb{R}^{H\times W\times C}$ into latent representations $\mathbf{X}\in\mathbb{R}^{h\times w\times c}$ using a pre-trained stable diffusion VAE encoder [29]. Following ViT [8], we embed each latent feature into non-overlapping patches $\mathbf{Z}\in\mathbb{R}^{t\times d}$, where $t=\frac{H}{p}\times\frac{W}{p}$ denotes the token count of each image, $p$ denotes the patch size, and apply random masking with ratios uniformly sampled from 0% to 90%. The masking ratios are carefully designed, low mask ratios are used to encourage the model to retain more complete information, while high mask ratios aim to enhance the model’s ability to abstract. Once the masking is complete, we discard the masked tokens and retain the unmasked tokens, arranging them into a sequence in their original order. As shown in Fig. 10, we demonstrate that even with 90% of the content masked, the model is still able to generate acceptable results.
$\small\mathbf{M}=\operatorname{Enc}(\mathbf{Z^{m}}\oplus\mathbf{T})$ (1)
where $\oplus$ denotes concatenation, $\operatorname{Enc}$ denotes Motion Encoder, $\mathbf{Z^{m}}$ denotes masked image patches.
TikTok [45] has demonstrated that image features can be represented as 1D sequence representations. Therefore, we designed a 1D learnable token sequence, which is adjustable in length and can become very compact. As shown in Fig. 4, we concatenate the learnable tokens with the image features and process them through ViT layers. Ultimately, latent motion $\mathbf{M}$ is output at the position corresponding to the learnable tokens. Here, the number of motion tokens and the token dimension size are consistent with the learnable tokens. Note that, unless otherwise specified, our latent motion is set to 1 token with a dimension of 512. Tab. 5 Tab. 6 varies number and dimension of tokens.
Diffusion Decoder. At the stage, our objective is to use the reference image as a reference for detailed information, with motion serving as the semantic condition, guiding a generative model to fill in missing details. Our decoder is a standard DiT [24] model. Specifically, we randomly select two sets of images from the data, obtaining reference image $\mathbf{Z_{r}}$ and target image feature $\mathbf{Z_{t}}$. We apply the same masking ratio to image features and obtain reference motion $\mathbf{M_{r}}$ and target motion $\mathbf{M_{t}}$ through the motion encoder. Following RectifiedFlow[20], We add noise to the unmasked target image feature Eq. 2, and concatenate noised target image feature $\mathbf{Z_{t}^{n}}$ with the $\mathbf{Z_{r}}$, $\mathbf{M_{r}}$ and $\mathbf{M_{t}}$ Eq. 3. Finally, the DiT decoder outputs the predicted noise velocity $\hat{\mathbf{V}}$. We refer to the module described above as the Transfer Block.
$\small\mathbf{Z_{t}^{n}}=(1-t)\mathbf{Z_{t}}+t\epsilon$ (2)
$\small\hat{\mathbf{V}}=\operatorname{Dec}(\mathbf{Z_{r}}\oplus\mathbf{Z_{t}^{n}}\oplus\mathbf{M_{r}}\oplus\mathbf{M_{t}})$ (3)
Where $\epsilon$ denotes Gaussian Noise sampled from standard normal distribution, $t\in[0,1]$ denotes the intensity of the noise, $\mathbf{Z_{t}^{n}}$ denotes noisy target image feature.
The above module enables the transfer from a reference image to a target image. We can extend this framework to transfer from the reference image to target video. Specifically, we randomly select one reference frame and a sequence of $f$ consecutive video frames. After applying Transfer Block to each frame, we introduce a Temporal Align Block. This is because we observed that treating each frame independently for denoising results in a lack of smooth transitions in the details. In Temporal Align Block, We flatten noised target image feature $\mathbf{Z^{n}_{t}}\in\mathbb{R}^{f\times t\times d}$ to $\mathbf{Z^{n}_{t}}\in\mathbb{R}^{b\times d}$, where $f$ denotes number of frames, $b=f\times t$. After transformation, we apply self-attention for temporal modeling, The complexity of this block is $o(f)$, which does not lead to significant computational overhead.
3.2 Motion Generator
In this module, our objective is to train a motion generator that can generate a corresponding motion sequence based on a previous motion sequence, a randomly selected reference motion, and an audio sequence. Our generator is a standard Diffusion Transformer (DiT). We first apply a 0–20% mask Eq. 4 to the previous image feature before extracting the reference motions $\mathbf{M_{p}^{m}}$ and extract target motions $\mathbf{M_{t}}$ without mask Eq. 5, we add noise to the target motions Eq. 6, these are concatenated along the time dimension and processed through self-attention layers. Following this, we use two cross-attention layers: one with the audio feature $\mathbf{A}$ and the other with the reference motion $\mathbf{M_{r}}$, note that the audio feature is embeded by Whisper [27]. In Fig. 9, we demonstrate that applying a mask to the previous image enhances the model’s generalization ability.
$\small\mathbf{M_{p}^{m}}=\operatorname{Enc}(\mathbf{Z^{m}_{p}}\oplus\mathbf{T})$ (4)
$\small\mathbf{M_{t}}=\operatorname{Enc}(\mathbf{Z_{t}}\oplus\mathbf{T})$ (5)
$\small\mathbf{M_{t}^{n}}=(1-t)\mathbf{M_{t}}+t\epsilon$ (6)
$\small\hat{\mathbf{V}}=\operatorname{G}\left(\mathbf{M_{p}^{m}},\mathbf{M_{t}^{n}},\mathbf{M_{r}},\mathbf{A}\right)$ (7)
3.3 Training Objective
In the first stage Eq. 8, the goal of the motion autoencoder is to reconstruct the target images based on the reference image and corresponding motion. In the second stage Eq. 9, the model is optimized to generate motion sequence based on audio condition. In both stage, the model predicts the noise velocity, and we optimize the mean squared distance between the predicted velocity $v_{\theta}$ and the true velocity $v$.
$\mathcal{L}_{1}=\mathbb{E}_{t,z_{r},z_{t},\epsilon}\left[\left\|(z_{t}-\epsilon)-v_{\theta}\left(z_{r},z_{t},t\right)\right\|^{2}\right]$ (8)
$\mathcal{L}_{2}=\mathbb{E}_{t,m_{r},m_{t},m_{p},\epsilon}\left[\left\|(m_{t}-\epsilon)-v_{\theta}\left(m_{r},m_{t},m_{p},t\right)\right\|^{2}\right]$ (9)
4 Experiment
4.1 Experimental Settings
Figure 6: Qualitative comparison with existing approaches. The results demonstrate that our model effectively generates complex, realistic, and natural motions across various audio conditions.
Table 1: Model Capacity of various methods.
Method Unsupervised Long-term Modeling Real-time Generation Generic Generation Pipeline
Aniportrait [41]
$\times$
$\times$
$\times$
$\times$
Hallo2 [6]
$\times$
$\times$
$\times$
$\times$
SadTalker [48]
$\times$
$\times$
Anitalker [21]
$\times$
Echomimic [4]
$\times$
$\times$
$\times$
$\times$
SeMo (Ours)
Table 2: Motion-based video reconstruction comparison of various methods on the HDTF, DH-FaceVid-1K, and MEAD dataset.
Method HDTF DH-FaceVid-1K MEAD
[…]
FOMM[30] 24.690 188.797 0.064 25.572 0.816 73.484 30.052 236.578 0.061 25.542 0.844 60.701 41.541 166.950 0.040 31.131 0.917 60.267
MRAA[31] 23.004 155.53 0.046 29.411 0.891 73.653 22.258 216.034 0.044 30.295 0.913 60.481 35.378 167.186 0.030 34.245 0.950 50.558
TPSMM[51] 26.556 138.88 0.041 29.51 0.901 58.478 24.697 215.362 0.042 29.625 0.917 54.784 34.358 136.651 0.028 34.805 0.956 69.224
LIA[39] 26.366 186.307 0.067 26.624 0.840 66.068 33.409 238.101 0.067 26.876 0.861 54.488 47.152 170.724 0.048 31.124 0.921 45.276
Liveportrait[12] 22.469 137.935 0.068 24.810 0.802 59.955 26.502 170.360 0.059 25.681 0.839 52.026 56.700 88.304 0.038 31.661 0.923 44.817
SeMo (Ours) 21.431 136.738 0.043 29.12 0.912 57.849 22.045 112.166 0.036 32.132 0.914 51.437 23.546 70.975 0.019 36.233 0.949 44.517
Datasets. For training the Motion AutoEncoder, we utilized three datasets: DH-FaceVid-1K [7], HDTF [50], and MEAD [36]. Videos were resized and center-cropped to 256×256 resolution, excluding those failing face detection. Training samples were equally drawn from all three datasets. For training the Motion Generator, we employed two datasets: HDTF and DH-FaceVid-1K. We removed samples with excessively noisy or unclear audio and encoded the audio using the pre-trained model Whisper [27].
Implementation Details. All two-stage experiments were trained on 8 NVIDIA A800 GPUs. In Stage 1, we first trained an image-to-image version of the Motion AutoEncoder. Subsequently, we utilized the pre-trained weights to further train it as an image-to-video version. In the main experiments, we maintained the motion representation as 1 token with dimension 512. In Stage 2, we trained the motion generator, generating 16 consecutive motion tokens in a single forward pass. A consistent learning rate of 1e-4 was applied across all experiments.
Evaluation Metrics. We evaluate our experiments very rigorously using various metrics. In the video reconstruction phase, we use FID [16], FVD [38], LPIPS [47], SSIM [40] and PSNR. In the audio-driven portrait video generation phase, we use FID, FVD, Sync-C and Sync-D.
Baseline. In the motion-based video reconstruction phase, we compared MRAA [31], FOMM [30], Lia [39], Liveportrait [12], TPSMM [51]. In the portrait video generation phase, we compared Aniportrait [41], Anitalker [21], Echomimic [4], Hallo2 [6] , SadTalker [48]
4.2 Comparisons.
Motion-based Video Reconstruction. Motion-based video reconstruction refers to an algorithm that generates target frames based on the reference image and motion. In the experiment, Tab. 2 we directly applied their pretrained weights in the face reconstruction task. We observed that our method outperformed theirs in terms of metrics. Additionally, our approach not only modeled the motion of the face but also modeled the motion in the surrounding areas of the face, making the entire video look more realistic.
Audio-driven Video Generation. Our generation framework achieves remarkably natural motion effects. As shown in Fig. 5 , when processing utterances like ”Yes!”, existing methods only produce limited lip movements, while our approach naturally integrates lip pursing with spontaneous eye blinks to deliver realistic contemplative expressions. Fig. 3 highlights our framework’s robustness across diverse facial poses and camera distances, generating substantial yet plausible motion variations. Notably, our method effectively captures authentic lighting interactions (e.g., dynamic shadow transitions during head movements), significantly enhancing visual realism. Quantitative evaluations in Tab. 4 Tab. 3 confirm our superiority over SOTA methods, with particularly significant improvements in FVD.
Tab. 1 demonstrates the strong capabilities of the model. The framework’s scalability stems from our compact latent motion representation, enabling long-term temporal modeling for holistic event understanding. Through RectifiedFlow’s reflow distillation, we achieve real-time synthesis with minimal sampling steps. As a prior-free and unsupervised architecture, our framework demonstrates strong cross-domain adaptability, enabling efficient application to novel scenarios without architectural modifications.
Table 3: Audio-driven video generation comparison of various methods on the DH-FaceVid-1K dataset.
Method DH-FaceVid-1K
[…]
AniPortrait [41] 44.807 484.861 0.168 19.103 0.671
Hallo2 [6] 61.803 698.668 0.155 20.016 0.658
SadTalker [48] 85.447 610.627 0.328 14.382 0.538
Anitalker [21] 56.322 556.607 0.289 15.064 0.572
Echomimic [4] 69.296 691.534 0.327 14.218 0.527
SeMo (Ours) 27.667 206.523 0.095 22.954 0.768
Table 4: Audio-driven video generation comparison of various methods on the HDTF dataset.
Method HDTF
[…]
AniPortrait [41] 33.139 422.644 0.138 20.042 0.672 5.817 10.881
Hallo2 [6] 27.447 276.564 0.105 21.692 0.713 6.284 8.499
SadTalker [48] 93.744 548.635 0.225 14.493 0.437 5.439 9.940
Anitalker [21] 41.034 430.082 0.208 17.434 0.588 5.927 8.724
Echomimic [4] 70.768 591.313 0.311 14.920 0.452 6.030 10.462
SeMo (Ours) 26.860 218.414 0.111 21.985 0.739 6.147 8.574
User Study. We randomly selected 10 images from HDTF and DH-FaceVid-1K, and generated 20 videos using 2 random audio clips. We invited 20 participants to choose the best model from 4 different perspectives. We significantly outperform competitors in realism and match the state-of-the-art (SOTA) in lip sync, smoothness, and overall quality.
4.3 Ablation Study
Dimension and Token Number. In Tab. 5 Tab. 6, we compared the impact of latent motion token dimensions and token numbers on model performance. For the token dimension, we fixed the number at 1 and varied its size, finding that even with a dimension as small as 32, the model could still generate reasonable images. However, smaller dimensions resulted in reduced detailed information, requiring the diffusion model to compensate, which introduced some randomness in the output, though the layout remained consistent. For the token number, we fixed the motion representation space at 512 floats and found that increasing the number of tokens improved performance. This is due to the increased parameters used by the model to analyze motion, enabling more accurate mapping from semantic tokens to low-level pixels.
Table 5: Ablation studies of the latent Motion’s channel dimension on the HDTF dataset.
Token Dimension HDTF
[…]
T1D32 29.877 237.652 0.068 26.916 0.824
T1D64 26.174 214.236 0.066 27.520 0.835
T1D128 24.467 192.702 0.053 27.587 0.837
T1D256 22.629 157.235 0.048 28.430 0.865
T1D512 21.431 136.738 0.043 29.12 0.871
T1D768 21.424 135.678 0.042 29.385 0.872
T1D1024 21.369 133.547 0.039 29.430 0.877
Table 6: Ablation studies of the latent Motion’s token length on the HDTF dataset.
Token Number HDTF
[…]
T1D512 21.431 136.738 0.043 29.120 0.871
T2D256 21.082 133.529 0.041 29.561 0.873
T4D128 20.561 130.061 0.040 29.992 0.880
T8D64 20.138 128.390 0.038 30.180 0.883
Figure 7: Unmasked Motion AutoEncoder training retains details in latent motion, while masking reduces them.
Mask Strategy. In Fig. 10 We compared the impact of using a mask strategy during the training of both the Motion AutoEncoder and Motion Generator. In training the Motion AutoEncoder, we applied a mask to both the reference and target image features, sampling the mask ratio from 0% to 90%. While reconstruction metrics showed similar results with or without the mask, we found that the version without the mask included more detailed motion information, aiding the diffusion decoder. However, the mask did not significantly affect reconstruction outcomes, and even with the mask, good results were achieved. When replacing the reference image with a different person, the version without the mask captured irrelevant details like background and clothing, while the masked version focused more on motion-related details from the reference image, aligning with our goal of abstracting motion while maintaining reconstruction quality. In training the Motion Generator, we applied a mask of 0-20% to reduce the generator’s reliance on detailed information from the previous motion and encourage it to extract details from the reference motion. Without the mask, the model struggled to generalize to other datasets, but with the mask, its generation capabilities improved and it became more adaptable to new datasets.
Table 7: Ablation studies of the Mask Strategy
Method HDTF
[…]
w/ mask 21.431 136.738 0.043 29.120 0.871
w/o mask 21.326 136.229 0.041 29.061 0.876
Figure 8: User study. The numbers in the figure indicate the win rate for each dimension.
Figure 9: Unmasked motion encoding causes overfitting to details, compromising cross-dataset generalization.
5 Discussion and Limitation
Figure 10: Motion encoding in SeMo requires minimal information, achieving good results even when 90% of the data is masked.
Figure 11: SeMo can be adapted to general video scenes.
Motion encoding requires minimal information. As demonstrated in the MAE [14], model needs very minimal to generate meaningful visual content. However, in the context of motion encoding, even less information is required. The core difference between motion encoding and image encoding lies in the fact that motion encoding has a reference image. This makes much of the detailed information in the target image unnecessary. Previous methods did not make good use of the information in the reference image. In the example Fig. 10, we observe that even 90% of the image is masked, the original video can still be reconstructed.
Transfer to General Scenarios. We investigated whether Motion AutoEncoder is applicable to general videos. In Fig. 11, We found that in single-subject motion scenarios, Motion AutoEncoder performs well in reconstruction, even effectively learning hand movements. However, in complex multi-subject motion scenarios, noticeable distortions occur. We reproduced IMF and train them on WebVid.
Abstract Resoning and Detail Generation. Current video generation models, like Sora, perform temporal reasoning and pixel generation simultaneously, neglecting redundancy in motion representation. Consider how humans think – we don’t fully visualize the next state, but rather abstract conceptual representations for efficient reasoning. Video generation models could adopt similar principles: by constructing a more compact informational space through abstraction, we enable efficient inference. Subsequently, the abstract information guides the generative model in pixel synthesis, making prediction more efficient and manageable.
Limitation. We found that the generated results have blurry teeth, and high-frequency areas like beards tend to flicker. We believe this is caused by the factor that: we resize data to 256x256 and compression into the latent space via SD VAE, there was significant information loss. However, our method is scalable, and we can easily extend it to higher resolutions, by maintaining the current approach.
6 Conclusion
In this work, we have proposed a fully self-supervised framework for portrait video generation, structured around three key steps: Abstraction, Reasoning, and Generation. First, we employ a masking strategy and a video reconstruction task to compress motion into a compact, abstract space (1D token). Second, within this space, we perform long-term modeling and efficient reasoning. Finally, the inferred motion serves as a conditioning signal to guide the diffusion model in generating fine-grained details. Furthermore, our findings indicate that motion can be further compressed, enabling a more compact and efficient abstract representation. Beyond portrait video generation, our pipeline is broadly applicable to general video generation tasks. In future research, we aim to explore its feasibility across a wider range of scenarios.

### v2

@@
-Semantic Latent Motion for Portrait Video Generation
+A Self-supervised Motion Representation for Portrait Video Generation
 Abstract
 Recent advancements in portrait video generation have been noteworthy.
-However, existing methods rely heavily on human priors and pre-trained generation models, which may introduce unrealistic motion and lead to inefficient inference.
+However, existing methods rely heavily on human priors and pre-trained generative models, Motion representations based on human priors may introduce unrealistic motion, while methods relying on pre-trained generative models often suffer from inefficient inference.
 To address these challenges, we propose Semantic Latent Motion (SeMo), a compact and expressive motion representation.
@@
 SeMo follows an effective three-step framework: Abstraction, Reasoning, and Generation.
-First, in the Abstraction step, we use a carefully designed Mask Motion Encoder to compress the subject’s motion state into a compact and abstract latent motion (1D token).
-Second, in the Reasoning step, long-term modeling and efficient reasoning are performed in this latent space to generate motion sequences.
-Finally, in the Generation step, the motion dynamics serve as conditional information to guide the generation model in synthesizing realistic transitions from reference frames to target frames.
-Thanks to the compact and descriptive nature of Semantic Latent Motion, our method enables real-time video generation with highly realistic motion.
+First, in the Abstraction step, we use a carefully designed Masked Motion Encoder, which leverages a self-supervised learning paradigm to compress the subject’s motion state into a compact and abstract latent motion (1D token).
+Second, in the Reasoning step, we efficiently generate motion sequences based on the driving audio signal.
+Finally, in the Generation step, the motion dynamics serve as conditional information to guide the motion decoder in synthesizing realistic transitions from reference frame to target video.
+Thanks to the compact and expressive nature of Semantic Latent Motion, our method achieves efficient motion representation and high-quality video generation.
 User studies demonstrate that our approach surpasses state-of-the-art models with an 81% win rate in realism.
 Extensive experiments further highlight its strong compression capability, reconstruction quality, and generative potential.
-Moreover, its fully self-supervised nature suggests promising applications in broader video generation tasks.
 1 Introduction
@@
 With the development of AI generation technologies, there is an increasing demand for more realistic animation effects and faster generation speeds.
-Figure 1: Semantic Latent Motion is a compact and expressive representation of the character’s current state.
-The motion dynamics in the semantic latent space serve as a powerful guiding condition, steering the generative model from the reference image to the target image.
-Leveraging the compactness of latent motion, we can model long sequences and perform more efficient resoning.
-Figure 2: Overview of SeMo.
-The proposed framework follows a three-step process (abstraction, reasoning, generation) implemented in two training stages.
-In the first stage, we train a Motion AutoEncoder for abstraction and generation.
-During the abstraction step, image features are subjected to random masking and compression to obtain a compact and abstract latent motion.
-In the generation step, latent motion is used as a condition to guide the diffusion decoder in generating the target image from the reference image.
-In the second stage, we train a Motion Generator for reasoning, which generates motion sequences from audio.
+Figure 1: Comparison of Motion Representation and Generation Methods.
+Motion representation: (a) Explicit motion representation (e.g., 3DMM, facial landmarks).
+(b) Implicit motion representation guided by explicit priors (e.g., implicit keypoints).
+(c) Our self-supervised implicit representation compresses motion into a compact 1D token without pretrained models or prior knowledge.
+Video generation approaches: (d) Adapter-modulated generation fine-tunes a pretrained video model via driving-signal adapters.
+(e) Identity-motion feature wrapping fuses identity features with target motion.
+(f) Our method leverages reference and target motion as conditions for efficient generation.
 Existing approaches mainly follow two paths.
-The first path integrates human prior knowledge of motion representation, such as sequences of 3DMM, pose, landmarks, etc.
-as expressions of motion [9, 4], combined with a generative renderer to produce the final video frames.
-The second path utilizes pre-trained video generation models, using their strong video generation capabilities and conditional control of cross-attention to learn the mapping from audio to video frames [6, 4, 41].
+In the first path (fig.
+1e), methods directly utilize explicit motion representations, or employ explicit priors to learn an implicit motion representation; these motion representations are then fused with identity features to generate video frames.
+In the second path (fig.
+1d), pre-trained video generation models are used, utilizing their powerful generative capacity and conditional cross-attention mechanisms to learn the mapping from the driving signal to video frames [6, 4, 42].
 Many methods combine both of these paths.
 However, in the first path, human prior knowledge relies on domain-specific heuristics, often failing to capture detailed motions, which results in human videos that lack naturalness and coherence.
-Moreover, this framework is not adaptable to multiple scenarios (for example, it cannot be used for general video generation, as it is impractical to extract pose sequences for all objects in the scene).
-In the second path, the complex spatiotemporal attention modeling and multi-step denoising in diffusion models prevent real-time generation and long-context modeling.
+Moreover, this framework is not adaptable to multiple scenarios (e.g., 3DMM is only applicable to facial scenarios and cannot be generalized to other types of scene).
+In the second path, the computationally intensive spatial-temporal attention operations - with their quadratic scaling in memory and computation - combined with multistep denoising in diffusion models create significant bottlenecks for video generation efficiency.
 We observe that these motion representations (3DMM, pose, landmarks) all express a high-level semantic:“ the positions and attributes of meaningful parts.” They are all abstractions of human knowledge about facial dynamics and are limited by design biases.
@@
 However, these representations are limited by human abstracting capabilities and cannot capture detailed motion.
-We ask: Can deep learning models abstract a semantic latent representation that better captures detailed face motions by self-supervised learning?
-We will introduce our motivation from the following two aspects.
-1) Representing video motion in a compact and expressive space.
-Modeling video motion presents significant challenges.
-Fortunately, facial videos exhibit more constrained motion ranges and patterns compared to general videos.
-Previous works have demonstrated that models can synthesize realistic images by extracting detailed information from reference images and semantic information from prior knowledge.
-Human-crafted priors (e.g., facial landmarks) represent compact and semantic motion representations abstracted through domain knowledge.
-If humans can achieve this, deep learning models could potentially learn superior representations through self-supervision – capturing more compact and detail-preserving motion characteristics.
-This motivates our exploration of learning semantic latent motion representations that surpass human-designed priors in expressiveness.
-Figure 3: Given an audio clip and a reference image, SeMo can generate lifelike video.
-From top to bottom, we highlight several key features of our model: a.
-The generated characters exhibit natural head movements and blinking while speaking.
-b.
-They display rich motion dynamics with substantial amplitude.
-c.
-Model can adapt to different poses.
-d.
-Model generalize well across various datasets.
-e.
-The generated videos feature remarkably natural lighting and shadow variations.
-2) Employing self-supervised semantic representations to guide generation.
+We investigate the fundamental question: How can we derive a motion representation that effectively captures fine-grained facial dynamics without relying on prior knowledge?
+To address this, we first present our motivation from two key perspectives.
+1) Video motion in a highly compact and semantic space.
+Prior works [34, 49, 40] have demonstrated that the content between frames in general videos contains substantial redundancy, video motion can be compressed into a very small latent space.
+For portrait videos in particular, the motion can be represented even more compactly, as the motion range of human subjects is inherently limited and the background typically remains static.
+Previous approaches leverage human-crafted semantic priors (such as 3DMM or landmarks) to represent motion through domain knowledge, and have validated the effectiveness of this strategy.
+This suggests that, by surpassing the abstraction capacity and compression limitations of human-designed priors, it may be possible to encode portrait video motion into an even more compact and semantic latent space.
+2) Self-supervised representations as guidance.
 Current generative models [5, 16, 1, 2, 44, 21, 50, 12] predominantly rely on human-abstracted conditions like text prompts or landmarks.
@@
 Recent works like RCG [17] and StoryDiffusion[56] have demonstrated that self-supervised latent features can serve as superior generation conditions compared to manual priors.
-Building on this insight, we propose using learned semantic latent motion as conditional signals for video generation, enabling more precise and detailed motion control.
-Motivated by these insights, we design a universal motion representation framework coupled with a simple yet effective video generation pipeline.
-Our approach consists of two key components: First, we train a Motion Autoencoder, where an encoder compresses each facial video frame into a compact latent motion representation (1D token) using a random masking strategy.
-The diffusion-based decoder then reconstructs the target frames using reference images as inputs and semantic latent motion as conditional signals.
-Once the Motion Autoencoder is trained, we extract latent motion representations for downstream tasks.
-For portrait video generation, we train a diffusion-based generator that synthesizes latent motion sequences directly from audio inputs.
-The compact latent space enables efficient long-term temporal modeling and preserves comprehensive event information.
-During inference, we first generate motion sequences from audio and reference images, then synthesize video frames using the Motion Autoencoder’s decoder.
-The rigorous reconstruction task encourages our latent motion representation to develop strong descriptive power, effectively capturing essential information within an extremely compact space.
-Leveraging this representation, our method achieves highly natural and vivid animations, outperforming state-of-the-art methods with an FID score of 26.8 (compared to the 27.4 baseline) and achieving an 81% win rate in realism in a user study.
-Beyond these results, we further observe that even when 90% of the video frame content (input to the motion encoder) is masked, our Motion AutoEncoder can still generate coherent and smooth videos.
-This reveals two critical insights: 1) High-level semantic information sufficiently guides video generation, 2) Video motion modeling requires less information proportion than image generation (contrasting 75% masking ratio in MAE [14]).
-Moreover, as a general framework, our Motion AutoEncoder demonstrates strong adaptability to generic videos containing limited subjects and simple motions, such as object rotations.
+Building on this insight, we propose using learned semantic latent motion as conditional signals for video generation.
+Motivated by these insights, we propose a universal motion representation framework paired with a simple yet effective video generation pipeline.
+Our approach comprises two key stages: In the first stage, we train a Motion Autoencoder in which the encoder compresses each video frame into a compact latent motion representation (1D token) using a random masking strategy.
+This encourages the model to abstract relevant semantic motion while discarding irrelevant details.
+The rectified flow-based decoder then leverages its powerful generative capabilities to fill in visual details, conditioned on the reference image and reference motion.
+In the second stage, we train a standard DiT model to directly generate motion sequences from audio and reference motions.
+Thanks to the highly compact nature of our motion representation, this process is extremely efficient.
+Our motion extraction and video generation pipeline is both simple and straightforward to generalize to other scenarios, requiring neither pretrained models nor hand-crafted human priors.
+Benefiting from the strong descriptive power of our semantic latent motion representation, our method produces highly natural and vivid animations, surpassing state-of-the-art approaches with an FID score of 26.8 (compared to a baseline of 27.4) and achieving an 81% win rate in realism according to user studies.
+In summary, our contributions to the community include:
+  1. 1.
+    We present a new representation framework that fuses the semantic descriptive capability of self-supervised representations with the fine-grained generative fidelity of a rectified-flow model, offering a new perspective on dynamic representation.
+  2. 2.
+    We show that the proposed representation encodes rich, complex motion cues in an compact latent space, leading to better reconstruction quality in video reconstruction tasks.
+  3. 3.
+    We validate the effectiveness of our approach in audio-driven portrait video generation.
+Owing to the self-supervised nature of SeMo, which eliminates the need for pretrained models or manually-designed priors, our method adapts well to a wide range of challenging scenarios while maintaining strong performance.
 2 Related Works
@@
 IMF [9] introduced a fully self-supervised approach that encodes reference and target frames into compact latent tokens.
-This method progressively reconstructs target frame pixels through a coarse-to-fine process that integrates reference image features at multiple granularity levels.
-These developments collectively demonstrate the evolution from explicit motion parameterization towards more implicit, learned representations that better preserve spatial-temporal coherence.
+The work most closely related to ours is IMF [9]appendix E illustrates the key differences between our approach and IMF.
 Portrait Video Generation.
@@
 Notably, VASA-1 [43] and Gaia [14] introduced hybrid architectures that encode both 3D prior knowledge and appearance features into a latent space, establishing a more universal and compact motion representation framework.
-Next, diffusion-based generators are employed to produce motion sequences conditioned on audio inputs [34, 19, 32, 49].
+Next, generative models are employed to produce motion sequences conditioned on audio inputs [33, 18, 31, 53].
 Recent innovations like EMO [33] and Hallo2 [6] have further enhanced visual quality through the integration of pre-trained video generation foundation models with specialized reference networks.
 AniTalker [19] presents a two-stage self-supervised learning paradigm.
-The overarching trend in this domain demonstrates a clear shift from hand-crafted priors towards data-driven approaches that enable models to capture finer details and enhance the naturalness of synthesized visuals.
 Self-supervised Representation as Guidance.
@@
 Meanwhile, Story Diffusion [56] obtains a continuous guidance signal by performing linear interpolation in the latent space of CLIP features, thereby controlling the video generation model to produce more coherent content.
+REPA [48] and VA-VAE [46] demonstrate the ability to improve generation and reconstruction by aligning latent features with strong self-supervised visual representations.
+Figure 2: Overview of SeMo.
+The proposed framework follows a three-step process (abstraction, reasoning, generation) implemented in two training stages.
+In the first stage, we train a Motion AutoEncoder for abstraction and generation.
+During the abstraction step, image features are subjected to random masking and compression to obtain a compact and abstract latent motion.
+In the generation step, latent motion is used as a condition to guide the motion decoder in generating the target video from the reference image.
+In the second stage, we train a Motion Generator for reasoning, which generates motion sequences from audio and reference motions.
 3 Method
-Figure 4: Illustration of Motion AutoEncoder
-Figure 5: An example of video reconstruction: In the example above, we demonstrate that SeMo can not only accurately model facial movements but also capture environmental changes.
-With a token size of 1, semantic latent motion is not only more compact than the motion in other works but also captures more comprehensive information.
 Given a reference image and its corresponding audio, our goal is to generate a dynamic video with synchronized motion and sound.
 To achieve this, we propose a two-stage self-supervised learning framework.
-In the first stage, the objective is to extract a compact, high-level semantic motion $\mathbf{M}\in\mathbb{R}^{l\times d}$ for each frame of the image.
-We first use learnable tokens and random masking to compress the motion into an abstract representation, and then utilize the reference image and the motion to reconstruct the target image.
-Once the Motion AutoEncoder is trained, it is frozen.
-In the second stage, we train a Motion Generator.
-For each frame, we first use the motion encoder to extract the motion and then generate the target motion based on the audio and the corresponding motion of the reference image.
+In the first stage, the objective is to extract a compact, high-level semantic motion $m\in\mathbb{R}^{l\times d}$ for each frame.
+We first use learnable tokens and random masking strategy to compress the motion into an abstract representation, and then utilize the reference image, reference motion and target motion to reconstruct the videos.
+In the second stage, we train a Motion Generator and the Motion AutoEncoder is frozen.
+For each frame, we first use the motion encoder to extract the motion and then generate motion sequence based on the audio and reference motion.
 The details of these two stages will be described below.
 3.1 Motion AutoEncoder
+Our core motivation is to leverage self-supervised learning to extract the most salient motion information, and then utilize the powerful distribution modeling ability of rectified flow to progressively complete the high-frequency details conditioned on the motion information in a coarse-to-fine manner.
 Motion Encoder.
 Our motion encoder design is motivated by the observation that facial movements - akin to 3D Morphable Models (3DMM) and keypoint trajectories - can be effectively represented in a highly compact space.
-We propose a novel Masked Motion Encoder architecture that integrates dual strategies of random masking and latent space compression to derive semantically rich latent motion.
-Specifically, we first compress input frames $\mathbf{I}\in\mathbb{R}^{H\times W\times C}$ into latent representations $\mathbf{X}\in\mathbb{R}^{h\times w\times c}$ using a pre-trained stable diffusion VAE encoder [29].
-Following ViT [8], we embed each latent feature into non-overlapping patches $\mathbf{Z}\in\mathbb{R}^{t\times d}$, where $t=\frac{H}{p}\times\frac{W}{p}$ denotes the token count of each image, $p$ denotes the patch size, and apply random masking with ratios uniformly sampled from 0% to 90%.
-The masking ratios are carefully designed, low mask ratios are used to encourage the model to retain more complete information, while high mask ratios aim to enhance the model’s ability to abstract.
+We propose a new Masked Motion Encoder architecture that integrates dual strategies of random masking and latent space compression to derive semantically rich latent motion.
+Specifically, we first compress input frames $I\in\mathbb{R}^{H\times W\times C}$ into latent representations $X\in\mathbb{R}^{h\times w\times c}$ using a pre-trained stable diffusion VAE encoder [27].
+Following ViT [8], we embed each latent into non-overlapping patches, and apply random masking with ratios uniformly sampled from 0% to 90%.
 Once the masking is complete, we discard the masked tokens and retain the unmasked tokens, arranging them into a sequence in their original order.
+The masking ratios are carefully designed: low mask ratios encourage the model to retain more complete information, while high mask ratios are intended to strengthen the model’s ability to ignore fine-grained details and focus on preserving the most salient motion information.
 As shown in fig.
-10, we demonstrate that even with 90% of the content masked, the model is still able to generate acceptable results.
-$\small\mathbf{M}=\operatorname{Enc}(\mathbf{Z^{m}}\oplus\mathbf{T})$ (1)
-where $\oplus$ denotes concatenation, $\operatorname{Enc}$ denotes Motion Encoder, $\mathbf{Z^{m}}$ denotes masked image patches.
+6, applying a masking strategy can greatly enhance the model’s generalization ability by ignoring unnecessary information.
+The mask strategy can enhance the model’s generalization ability fig.
+10 and also accelerate convergence during the second-stage training fig.
+7.
+More analysis is provided in fig.
+7.
 TikTok [47] has demonstrated that image features can be represented as 1D sequence representations.
-Therefore, we designed a 1D learnable token sequence, which is adjustable in length and can become very compact.
+Therefore, we designed a 1D learnable token $y\in\mathbb{R}^{l\times d}$ where $l$ denotes the number of tokens and $d$ denotes the features dimension.
+we concatenate the learnable token with the image features and process them through ViT layers.
+Ultimately, latent motion $M$ is output at the position corresponding to the learnable tokens.
+Here, the number of motion tokens and the token dimension size are consistent with the learnable tokens.
+Each image is encoded independently; we refer to the image features and motion associated with the reference image as $x_{\text{ref}}$,$m_{\text{ref}}$, and those associated with the target image as $x_{\text{tar}}$,$m_{\text{tar}}$.
+The encoding process can be formulated as follows: $m_{\text{ref}}=\operatorname{Enc}_{\theta}(x_{\text{ref}})$ , $m_{\text{tar}}=\operatorname{Enc}_{\theta}(x_{\text{tar}})$, where $\theta$ is the parameter of the motion encoder.
+Figure 3: Motion Decoder.
+The Motion Control Module uses the reference motion $m_{\text{ref}}$ and target motion $m_{\text{tar}}$ as control signals.
+The goal of this module is to transfer from the reference image features $x_{\text{ref}}$ to the target image features $x_{\text{tar}}$.
+After reshaping, the target image features $x_{\text{tar}}$ are processed with self-attention along the temporal dimension, enabling alignment of high-frequency information.
+Motion Decoder.
+At the decoding stage, our objective is to use the reference image for appearance information, with motion serving as the semantic condition, guiding a generative model to fill in missing details.
+Our decoder consists of two modules: the Motion Control Module and the Temporal Align Module (MCM and TAM for short).
+In the Motion Control Module, we concatenate the reference image features $x_{\text{ref}}$, the noised target image features $\bar{x}_{\text{tar}}$, the reference motion $m_{\text{ref}}$, and the target motion $m_{\text{tar}}$, then process them through a self-attention layer followed by a feed-forward network (FFN).
+The noising process follows the schedule in rectified flow, which can be formulated as: $\bar{x}=(1-t)x+t\epsilon$, where $t\in[0,1]$ is the noise level, and $\epsilon$ denotes Gaussian Noise sampled from standard normal distribution.
+$z_{\text{tar}}=\operatorname{MCM}_{\psi_{1}}(x_{\text{ref}},\bar{x}_{\text{tar}},m_{\text{ref}},m_{\text{tar}})$.
+Since denoising process in Motion Control Module is performed on each target frame, there could be inconsistencies in the high-frequency details across target frames.
+Therefore, in the Temporal Align Module, we concatenate the target image features $\{z_{\text{tar}}^{f}\}_{f=1}^{k}$ along the temporal dimension and apply self-attention to align their high-frequency information, where $k$ is the number of frames.
+$\{x_{\text{tar}}^{f}\}_{f=1}^{k}=\operatorname{TAM}_{\psi_{2}}(\{z_{\text{tar}}^{f}\}_{f=1}^{k})$.
+In all, the decoding process can be formulated as follows: $\{x_{\text{tar}}^{f}\}_{f=1}^{k}=\operatorname{Dec}_{\psi}(x_{\text{ref}},\{\bar{x}_{\text{tar}}^{f}\}_{f=1}^{k},m_{\text{ref}},\{m_{\text{tar}}^{f}\}_{f=1}^{k})$, where $\psi$ is the parameter of the motion decoder.
+3.2 Motion Generator
+Let $p$ be the number of previous frames input.
+In this module, our objective is to train a motion generator that can generate a corresponding motion sequence based on a previous motion $\{m_{\text{pre}}^{f}\}_{f=1}^{p}$, noised target motion $\{\bar{m}_{\text{tar}}^{f}\}_{f=1}^{k}$ and an audio sequence $\{a^{f}\}_{f=1}^{k}$.
+Note that $\{m_{\text{pre}}^{f}\}_{f=1}^{p}$ and $\{m_{\text{tar}}^{f}\}_{f=1}^{k}$ is extracted by motion encoder from corresponding image features without mask and the audio sequence $\{a^{f}\}_{f=1}^{k}$ is encoded by a pretrained model Whisper [25].
+Specifically, we concatenate the noised target motions $\{{\bar{m}}_{\text{tar}}^{f}\}_{f=1}^{k}$ with previous motion $\{m_{\text{pre}}^{f}\}_{f=1}^{p}$ for self-attention modeling, while cross-attention layers establish motion-audio alignment.
 As shown in fig.
-4, we concatenate the learnable tokens with the image features and process them through ViT layers.
-Ultimately, latent motion $\mathbf{M}$ is output at the position corresponding to the learnable tokens.
-Here, the number of motion tokens and the token dimension size are consistent with the learnable tokens.
-Note that, unless otherwise specified, our latent motion is set to 1 token with a dimension of 512.
-Tab.
-5 Tab.
-6 varies number and dimension of tokens.
-Diffusion Decoder.
-At the stage, our objective is to use the reference image as a reference for detailed information, with motion serving as the semantic condition, guiding a generative model to fill in missing details.
-Our decoder is a standard DiT [24] model.
-Specifically, we randomly select two sets of images from the data, obtaining reference image $\mathbf{Z_{r}}$ and target image feature $\mathbf{Z_{t}}$.
-We apply the same masking ratio to image features and obtain reference motion $\mathbf{M_{r}}$ and target motion $\mathbf{M_{t}}$ through the motion encoder.
-Following RectifiedFlow[20], We add noise to the unmasked target image feature Eq.
-2, and concatenate noised target image feature $\mathbf{Z_{t}^{n}}$ with the $\mathbf{Z_{r}}$, $\mathbf{M_{r}}$ and $\mathbf{M_{t}}$ Eq.
-3.
-Finally, the DiT decoder outputs the predicted noise velocity $\hat{\mathbf{V}}$.
-We refer to the module described above as the Transfer Block.
-$\small\mathbf{Z_{t}^{n}}=(1-t)\mathbf{Z_{t}}+t\epsilon$ (2)
-$\small\hat{\mathbf{V}}=\operatorname{Dec}(\mathbf{Z_{r}}\oplus\mathbf{Z_{t}^{n}}\oplus\mathbf{M_{r}}\oplus\mathbf{M_{t}})$ (3)
-Where $\epsilon$ denotes Gaussian Noise sampled from standard normal distribution, $t\in[0,1]$ denotes the intensity of the noise, $\mathbf{Z_{t}^{n}}$ denotes noisy target image feature.
-The above module enables the transfer from a reference image to a target image.
-We can extend this framework to transfer from the reference image to target video.
-Specifically, we randomly select one reference frame and a sequence of $f$ consecutive video frames.
-After applying Transfer Block to each frame, we introduce a Temporal Align Block.
-This is because we observed that treating each frame independently for denoising results in a lack of smooth transitions in the details.
-In Temporal Align Block, We flatten noised target image feature $\mathbf{Z^{n}_{t}}\in\mathbb{R}^{f\times t\times d}$ to $\mathbf{Z^{n}_{t}}\in\mathbb{R}^{b\times d}$, where $f$ denotes number of frames, $b=f\times t$.
-After transformation, we apply self-attention for temporal modeling, The complexity of this block is $o(f)$, which does not lead to significant computational overhead.
-3.2 Motion Generator
-In this module, our objective is to train a motion generator that can generate a corresponding motion sequence based on a previous motion sequence, a randomly selected reference motion, and an audio sequence.
-Our generator is a standard Diffusion Transformer (DiT).
-We first apply a 0–20% mask Eq.
-4 to the previous image feature before extracting the reference motions $\mathbf{M_{p}^{m}}$ and extract target motions $\mathbf{M_{t}}$ without mask Eq.
-5, we add noise to the target motions Eq.
-6, these are concatenated along the time dimension and processed through self-attention layers.
-Following this, we use two cross-attention layers: one with the audio feature $\mathbf{A}$ and the other with the reference motion $\mathbf{M_{r}}$, note that the audio feature is embeded by Whisper [27].
-In Fig.
-9, we demonstrate that applying a mask to the previous image enhances the model’s generalization ability.
-$\small\mathbf{M_{p}^{m}}=\operatorname{Enc}(\mathbf{Z^{m}_{p}}\oplus\mathbf{T})$ (4)
-$\small\mathbf{M_{t}}=\operatorname{Enc}(\mathbf{Z_{t}}\oplus\mathbf{T})$ (5)
-$\small\mathbf{M_{t}^{n}}=(1-t)\mathbf{M_{t}}+t\epsilon$ (6)
-$\small\hat{\mathbf{V}}=\operatorname{G}\left(\mathbf{M_{p}^{m}},\mathbf{M_{t}^{n}},\mathbf{M_{r}},\mathbf{A}\right)$ (7)
+7, the masking strategy improve the distribution of motion latents, which accelerates the convergence as well as the performance of the second-stage training.
+The motion generation process can be formulated as follows: $\{m_{\text{tar}}^{f}\}_{f=1}^{k}=\operatorname{Gen}_{\phi}(\{\bar{m}_{\text{tar}}^{f}\}_{f=1}^{k},\{m_{\text{pre}}^{f}\}_{f=1}^{p},\{a^{f}\}_{f=1}^{k})$, where $\phi$ is the parameter of the motion generator.
 3.3 Training Objective
-In the first stage Eq.
-8, the goal of the motion autoencoder is to reconstruct the target images based on the reference image and corresponding motion.
-In the second stage Eq.
-9, the model is optimized to generate motion sequence based on audio condition.
-In both stage, the model predicts the noise velocity, and we optimize the mean squared distance between the predicted velocity $v_{\theta}$ and the true velocity $v$.
-$\mathcal{L}_{1}=\mathbb{E}_{t,z_{r},z_{t},\epsilon}\left[\left\|(z_{t}-\epsilon)-v_{\theta}\left(z_{r},z_{t},t\right)\right\|^{2}\right]$ (8)
-$\mathcal{L}_{2}=\mathbb{E}_{t,m_{r},m_{t},m_{p},\epsilon}\left[\left\|(m_{t}-\epsilon)-v_{\theta}\left(m_{r},m_{t},m_{p},t\right)\right\|^{2}\right]$ (9)
+In the first stage, the goal of the motion autoencoder is to reconstruct the target video based on the reference image and corresponding motion.
+In the second stage, the model is optimized to generate motion sequence based on audio condition.
+Both stage utilize rectified flow as the generative framework.
+Let $v_{\psi}$ and $u_{\phi}$ denote the velocity respectively, we optimize the distance between the predicted velocity $v_{\theta}$ and the true velocity $v$.
+$\min_{\theta,\psi}\mathcal{L}_{\text{first stage}}=\min_{\theta,\psi}\mathbb{E}_{t,x_{\text{ref}},x_{\text{tar}},\epsilon}\left[\left\|(x_{\text{tar}}-\epsilon)-v_{\psi}\left(x_{\text{ref}},\bar{x}_{\text{tar}},\operatorname{Enc}_{\theta}(x_{\text{ref}}),\operatorname{Enc}_{\theta}(x_{\text{tar}}),t\right)\right\|^{2}\right]$ (1)
+$\min_{\phi}\mathcal{L}_{\text{second stage}}=\min_{\phi}\mathbb{E}_{{t},a,m_{\text{tar}},m_{\text{pre}},\epsilon}\left[\left\|(m_{\text{tar}}-\epsilon)-u_{\phi}\left(a,\bar{m}_{\text{tar}},m_{\text{pre}},t\right)\right\|^{2}\right]$ (2)
 4 Experiment
+Our experiments are mainly divided into three parts.
+First, we verify the effectiveness of the representations on the video reconstruction task section 4.2.1.
+Then, we evaluate the applicability of the representations to downstream tasks by testing on audio-driven portrait video generation section 4.2.2.
+Finally, we conduct extensive ablation studies to demonstrate the necessity of our experimental strategies section 4.3.
 4.1 Experimental Settings
-Figure 6: Qualitative comparison with existing approaches. The results demonstrate that our model effectively generates complex, realistic, and natural motions across various audio conditions.
-Table 1: Model Capacity of various methods.
-Method Unsupervised Long-term Modeling Real-time Generation Generic Generation Pipeline
-Aniportrait [41]
-$\times$
-$\times$
-$\times$
-$\times$
-Hallo2 [6]
-$\times$
-$\times$
-$\times$
-$\times$
-SadTalker [48]
-$\times$
-$\times$
-Anitalker [21]
-$\times$
-Echomimic [4]
-$\times$
-$\times$
-$\times$
-$\times$
-SeMo (Ours)
-Table 2: Motion-based video reconstruction comparison of various methods on the HDTF, DH-FaceVid-1K, and MEAD dataset.
+Figure 4: Results of video reconstruction. We show video reconstruction results on datasets DH-FaceVid-1K and HDTF.
+Figure 5: Result of Audio-driven Video Generation.
+Our method adapts to various challenging scenarios.
+Zoom in for a better view.
+(a) Close view: the subject is very close to the camera; (b) Distant view: the subject is far from the camera, nearly half-body is shown; (c) Edge view: the face appears at the edge of the frame; (d) Partial face: only part of the face is visible; (e) Head turn: the subject turns their head.
+Table 1: Motion-based video reconstruction comparison of various methods on the HDTF, DH-FaceVid-1K, and MEAD dataset. Red indicates the best result, blue the second best, and green the third best.
 Method HDTF DH-FaceVid-1K MEAD
 […]
-FOMM[30] 24.690 188.797 0.064 25.572 0.816 73.484 30.052 236.578 0.061 25.542 0.844 60.701 41.541 166.950 0.040 31.131 0.917 60.267
-MRAA[31] 23.004 155.53 0.046 29.411 0.891 73.653 22.258 216.034 0.044 30.295 0.913 60.481 35.378 167.186 0.030 34.245 0.950 50.558
-TPSMM[51] 26.556 138.88 0.041 29.51 0.901 58.478 24.697 215.362 0.042 29.625 0.917 54.784 34.358 136.651 0.028 34.805 0.956 69.224
-LIA[39] 26.366 186.307 0.067 26.624 0.840 66.068 33.409 238.101 0.067 26.876 0.861 54.488 47.152 170.724 0.048 31.124 0.921 45.276
-Liveportrait[12] 22.469 137.935 0.068 24.810 0.802 59.955 26.502 170.360 0.059 25.681 0.839 52.026 56.700 88.304 0.038 31.661 0.923 44.817
-SeMo (Ours) 21.431 136.738 0.043 29.12 0.912 57.849 22.045 112.166 0.036 32.132 0.914 51.437 23.546 70.975 0.019 36.233 0.949 44.517
+FOMM[29] 24.690 188.797 0.064 25.572 0.816 30.052 236.578 0.061 25.542 0.844 41.541 166.950 0.040 31.131 0.917
+MRAA[30] 23.004 155.53 0.046 29.411 0.891 22.258 216.034 0.044 30.295 0.913 35.378 167.186 0.030 34.245 0.950
+TPSMM[55] 26.556 138.88 0.041 29.51 0.901 24.697 215.362 0.042 29.625 0.917 34.358 136.651 0.028 34.805 0.956
+LIA[39] 26.366 186.307 0.067 26.624 0.840 33.409 238.101 0.067 26.876 0.861 47.152 170.724 0.048 31.124 0.921
+Liveportrait[11] 22.469 137.935 0.068 24.810 0.802 26.502 170.360 0.059 25.681 0.839 56.700 88.304 0.038 31.661 0.923
+SeMo (Ours) 21.431 136.738 0.043 29.12 0.912 22.045 112.166 0.036 32.132 0.914 23.546 70.975 0.019 36.233 0.949
 Datasets.
@@
 All two-stage experiments were trained on 8 NVIDIA A800 GPUs.
-In Stage 1, we first trained an image-to-image version of the Motion AutoEncoder.
-Subsequently, we utilized the pre-trained weights to further train it as an image-to-video version.
-In the main experiments, we maintained the motion representation as 1 token with dimension 512.
-In Stage 2, we trained the motion generator, generating 16 consecutive motion tokens in a single forward pass.
-A consistent learning rate of 1e-4 was applied across all experiments.
+We set the dimension of the semantic latent motion to 1 token with 512 channels throughout all experiments to balance reconstruction accuracy and generation quality.
+In the first stage, the motion autoencoder generates a sequence of $k=16$ frames in a single forward pass.
+In the second stage, the motion generator also produces $k=16$ motion tokens per forward pass aided with $p=8$ previous motion tokens.
+All models are trained under similar settings: a base learning rate of $10^{-4}$ per 64 batch size and the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\text{decay}=0.05$.
 Evaluation Metrics.
@@
 In the motion-based video reconstruction phase, we compared MRAA [30], FOMM [29], Lia [39], Liveportrait [11], TPSMM [55].
-In the portrait video generation phase, we compared Aniportrait [41], Anitalker [21], Echomimic [4], Hallo2 [6] , SadTalker [48]
-4.2 Comparisons.
-Motion-based Video Reconstruction.
+In the portrait video generation phase, we compared Aniportrait [42], Anitalker [19], Echomimic [4], Hallo2 [6], SadTalker [52].
+4.2 Comparison
+4.2.1 Motion-based Video Reconstruction.
 Motion-based video reconstruction refers to an algorithm that generates target frames based on the reference image and motion.
-In the experiment, Tab.
-2 we directly applied their pretrained weights in the face reconstruction task.
+In the experiment, table 1 we directly applied their pretrained weights in the face reconstruction task.
 We observed that our method outperformed theirs in terms of metrics.
-Additionally, our approach not only modeled the motion of the face but also modeled the motion in the surrounding areas of the face, making the entire video look more realistic.
-Audio-driven Video Generation.
-Our generation framework achieves remarkably natural motion effects.
-As shown in Fig.
-5 , when processing utterances like ”Yes!”, existing methods only produce limited lip movements, while our approach naturally integrates lip pursing with spontaneous eye blinks to deliver realistic contemplative expressions.
-Fig.
-3 highlights our framework’s robustness across diverse facial poses and camera distances, generating substantial yet plausible motion variations.
-Notably, our method effectively captures authentic lighting interactions (e.g., dynamic shadow transitions during head movements), significantly enhancing visual realism.
-Quantitative evaluations in Tab.
-4 Tab.
-3 confirm our superiority over SOTA methods, with particularly significant improvements in FVD.
-Tab.
-1 demonstrates the strong capabilities of the model.
-The framework’s scalability stems from our compact latent motion representation, enabling long-term temporal modeling for holistic event understanding.
-Through RectifiedFlow’s reflow distillation, we achieve real-time synthesis with minimal sampling steps.
-As a prior-free and unsupervised architecture, our framework demonstrates strong cross-domain adaptability, enabling efficient application to novel scenarios without architectural modifications.
-Table 3: Audio-driven video generation comparison of various methods on the DH-FaceVid-1K dataset.
+Additionally, Since the training task involves reconstructing entire videos through self-supervised learning rather than focusing solely on the face, in fig.
+4,we observe that the model is able to capture more detailed and complex motions, such as those of earrings and hairstyles.
+Furthermore, thanks to the powerful generative capability of the motion decoder, the model can automatically complete missing details caused by head movements.
+Methods based on explicit representations [29, 30, 55] perform poorly, especially in cases involving large or rapid motions.
+They often fail to capture fine-grained, high-frequency changes.
+Moreover, these explicit representations are typically designed for the face only, neglecting global motion information, which further limits their performance and robustness.
+As a result, such methods tend to generalize poorly and cannot be effectively transferred to other scenarios.
+We further provide extensive insights on the mask section D.1 and scene transferability section D.4.
+Figure 6: (a) Comparison of efficiency and quality among audio-driven video generation methods.
+Bits per Pixel is calculated as $\frac{S\cdot D}{256^{2}}$, where $S$ is the representation size and D is the bit width of each element, as proposed by  [28].
+(b) User study on audio-driven video generation.
+The detail procedure of user study can be found in appendix C.
+Table 2: Video generation comparison on DH-FaceVid-1K and HDTF.
 Method DH-FaceVid-1K
@@
 SeMo (Ours) 27.667 206.523 0.095 22.954 0.768
-Table 4: Audio-driven video generation comparison of various methods on the HDTF dataset.
 Method HDTF
@@
 SeMo (Ours) 26.860 218.414 0.111 21.985 0.739 6.147 8.574
-User Study.
-We randomly selected 10 images from HDTF and DH-FaceVid-1K, and generated 20 videos using 2 random audio clips.
-We invited 20 participants to choose the best model from 4 different perspectives.
-We significantly outperform competitors in realism and match the state-of-the-art (SOTA) in lip sync, smoothness, and overall quality.
+4.2.2 Audio-driven Video Generation.
+The quantitative metrics for audio-driven video generation are presented in table 2, while the qualitative (visualization) results are shown in fig.
+5.
+The Sync-C and Sync-D metric was not evaluated on the DH-FaceVid-1K dataset because the models used for evaluation were not trained on Chinese datasets.
+Our method not only achieves highly realistic visuals, but also excels in robustness across diverse scenarios.
+As shown in fig.
+5, it effectively handles challenging cases where others often fail.
+Competing methods lack eyeball detail in close-ups, produce unnatural head turns, and frequently fail or generate unrealistic content in partial-face situations.
+These limitations point to a common issue: prior-based methods [42, 6, 7, 4] require strict assumptions about scene conditions, making such methods heavily scenario-dependent.
+Typically, they rely on complex data pre-processing pipelines to tightly crop and center the face within the frame.
+In contrast, our self-supervised approach removes the dependency on prior knowledge, enabling robust performance in unconstrained settings.
+As shown in fig.
+6(b), our user study results demonstrate that our method achieves an impressive 81 % win rate for perceived realism.
+This indicates that our approach captures more nuanced facial motions, while others often produce noticeable rigid-body translations—a subtle difference easily perceived by human evaluators but not always captured by standard quantitative metrics.In fig.
+6(a), we compare the encoding efficiency of each method, which is calculated as the size of the latent representation per frame in the latent space divided by the number of pixels per frame.
+More visual results can be found in section D.3 and video materials.
 4.3 Ablation Study
-Dimension and Token Number.
-In Tab.
-5 Tab.
-6, we compared the impact of latent motion token dimensions and token numbers on model performance.
-For the token dimension, we fixed the number at 1 and varied its size, finding that even with a dimension as small as 32, the model could still generate reasonable images.
-However, smaller dimensions resulted in reduced detailed information, requiring the diffusion model to compensate, which introduced some randomness in the output, though the layout remained consistent.
-For the token number, we fixed the motion representation space at 512 floats and found that increasing the number of tokens improved performance.
-This is due to the increased parameters used by the model to analyze motion, enabling more accurate mapping from semantic tokens to low-level pixels.
-Table 5: Ablation studies of the latent Motion’s channel dimension on the HDTF dataset.
-Token Dimension HDTF
+Table 3: Ablation studies on the number of motion tokens and the token channel dimension, where T denotes the number of tokens and D denotes the channel dimension.
+Token HDTF
 […]
@@
 T1D1024 21.369 133.547 0.039 29.430 0.877
-Table 6: Ablation studies of the latent Motion’s token length on the HDTF dataset.
-Token Number HDTF
 […]
-T1D512 21.431 136.738 0.043 29.120 0.871
-T2D256 21.082 133.529 0.041 29.561 0.873
 T4D128 20.561 130.061 0.040 29.992 0.880
 T8D64 20.138 128.390 0.038 30.180 0.883
-Figure 7: Unmasked Motion AutoEncoder training retains details in latent motion, while masking reduces them.
-Mask Strategy.
+Dimension and Token Number.
+In table 3, we compared the impact of latent motion token dimensions and token numbers on model performance.
+For the token dimension, we fixed the number at 1 and varied its size, finding that even with a dimension as small as 32, the model could still generate reasonable images.
+However, smaller dimensions resulted in reduced detailed information, requiring the motion decoder to compensate, which introduced some randomness in the output, though the layout remained consistent.
+For the token number, we fixed the motion representation space at 512 floats and found that increasing the number of tokens improved performance.
+This is due to the increased parameters used by the model to analyze motion, enabling more accurate mapping from semantic tokens to low-level pixels.
+Figure 7: (a) Training loss of Stage 2 with and without mask. (b) UMap [22] visualization of latent motion. (c) Attention Map between latent motion and reference image.
+Masking.
 In  fig.
-10 We compared the impact of using a mask strategy during the training of both the Motion AutoEncoder and Motion Generator.
-In training the Motion AutoEncoder, we applied a mask to both the reference and target image features, sampling the mask ratio from 0% to 90%.
-While reconstruction metrics showed similar results with or without the mask, we found that the version without the mask included more detailed motion information, aiding the diffusion decoder.
-However, the mask did not significantly affect reconstruction outcomes, and even with the mask, good results were achieved.
-When replacing the reference image with a different person, the version without the mask captured irrelevant details like background and clothing, while the masked version focused more on motion-related details from the reference image, aligning with our goal of abstracting motion while maintaining reconstruction quality.
-In training the Motion Generator, we applied a mask of 0-20% to reduce the generator’s reliance on detailed information from the previous motion and encourage it to extract details from the reference motion.
-Without the mask, the model struggled to generalize to other datasets, but with the mask, its generation capabilities improved and it became more adaptable to new datasets.
-Table 7: Ablation studies of the Mask Strategy
-Method HDTF
-[…]
-w/ mask 21.431 136.738 0.043 29.120 0.871
-w/o mask 21.326 136.229 0.041 29.061 0.876
-Figure 8: User study. The numbers in the figure indicate the win rate for each dimension.
-Figure 9: Unmasked motion encoding causes overfitting to details, compromising cross-dataset generalization.
-5 Discussion and Limitation
-Figure 10: Motion encoding in SeMo requires minimal information, achieving good results even when 90% of the data is masked.
-Figure 11: SeMo can be adapted to general video scenes.
-Motion encoding requires minimal information.
-As demonstrated in the MAE [14], model needs very minimal to generate meaningful visual content.
-However, in the context of motion encoding, even less information is required.
-The core difference between motion encoding and image encoding lies in the fact that motion encoding has a reference image.
-This makes much of the detailed information in the target image unnecessary.
-Previous methods did not make good use of the information in the reference image.
-In the example Fig.
-10, we observe that even 90% of the image is masked, the original video can still be reconstructed.
-Transfer to General Scenarios.
-We investigated whether Motion AutoEncoder is applicable to general videos.
+7(a), we find that masked training accelerates Stage 2 convergence, lowers loss, and improves output quality.
+We infer that masking forces the model to prioritize high-level motion semantics over local details, making generation more predictable.
 In  fig.
-11, We found that in single-subject motion scenarios, Motion AutoEncoder performs well in reconstruction, even effectively learning hand movements.
-However, in complex multi-subject motion scenarios, noticeable distortions occur.
-We reproduced IMF and train them on WebVid.
-Abstract Resoning and Detail Generation.
-Current video generation models, like Sora, perform temporal reasoning and pixel generation simultaneously, neglecting redundancy in motion representation.
-Consider how humans think – we don’t fully visualize the next state, but rather abstract conceptual representations for efficient reasoning.
-Video generation models could adopt similar principles: by constructing a more compact informational space through abstraction, we enable efficient inference.
-Subsequently, the abstract information guides the generative model in pixel synthesis, making prediction more efficient and manageable.
+7(b), we observe that masking encourages discrete latent motion features, aligning with findings in [46] where discrete distributions enhance generation.
+In  fig.
+7(d), We observe that, guided by the latent motion, the model mainly focuses on regions such as the eyes and mouth, where the movements are more significant.
+In  fig.
+9(c), we observe that using a masking strategy can partially separate identity and motion: replacing the reference video with another identity shows less identity leakage when masking is applied.
+We attribute this effect to the randomly sampled mask ratios, which encourage the model to abstract at high ratios and retain some facial details at lower ones.
+5 Conclusion and Limitations
+Portrait video generation has a wide range of applications, but it faces challenges in both generation efficiency and quality.
+Moreover, existing methods often rely on prior knowledge, which limits their applicability to fixed facial scenarios.
+In this work, we address these issues by proposing a self-supervised motion representation framework SeMo, which is clear and straightforward, consisting of three steps: Abstraction, Reasoning, and Generation.
+Thanks to the compact and expressive nature of SeMo, we demonstrate that SeMo significantly improves generation efficiency and quality.
+Beyond portrait video generation, our pipeline is broadly applicable to general video generation tasks.
+In future work, we plan to further explore its feasibility and effectiveness across a wider range of scenarios.
 Limitation.
-We found that the generated results have blurry teeth, and high-frequency areas like beards tend to flicker.
-We believe this is caused by the factor that: we resize data to 256x256 and compression into the latent space via SD VAE, there was significant information loss.
-However, our method is scalable, and we can easily extend it to higher resolutions, by maintaining the current approach.
-6 Conclusion
-In this work, we have proposed a fully self-supervised framework for portrait video generation, structured around three key steps: Abstraction, Reasoning, and Generation.
-First, we employ a masking strategy and a video reconstruction task to compress motion into a compact, abstract space (1D token).
-Second, within this space, we perform long-term modeling and efficient reasoning.
-Finally, the inferred motion serves as a conditioning signal to guide the diffusion model in generating fine-grained details.
-Furthermore, our findings indicate that motion can be further compressed, enabling a more compact and efficient abstract representation.
-Beyond portrait video generation, our pipeline is broadly applicable to general video generation tasks.
-In future research, we aim to explore its feasibility across a wider range of scenarios.
+Due to the need for efficient generation and better user accessibility, we limited our experiments to a resolution of 256×256.
+However, this downsampling, together with compression by the SD VAE, causes some information loss, resulting in blurry teeth and flickering in high-frequency regions such as beards.
+In future work, we plan to explore higher resolutions to further improve the visual quality of the generated results.
+Appendix A Motivation
+Our motivation stems from the observation that video reasoning is inherently high-dimensional and complex; predicting every pixel is highly uncertain, and humans do not explicitly forecast each pixel in their brains.
+Instead, we tend to abstract concepts from what we see and perform efficient reasoning in this conceptual space, which not only improves the reliability but also the efficiency of predictions.
+Inspired by this, we apply a similar principle to portrait video generation.
+Rather than directly predicting pixels, the model first abstracts each frame into high-level concepts, enabling more effective prediction within this compact space.
+Leveraging the strong generative and conditional capabilities of diffusion models, we then progressively refine predictions from high-level semantic concepts to final pixel-level outputs in a coarse-to-fine manner.
+As illustrated in the 8, given a reference image, reference motion, and target motion, the model can easily generate the target frame at the pixel level.
+To this end, we combine self-supervised learning and diffusion-based generative modeling in a three-step framework: Abstraction, reasoning, and generation.
+Figure 8: Motivation.
+Appendix B Implementation Details
+As described in Sec.
+3.1, SeMo adopts an EncoderDecoder architecture.
+Specifically, In the motion encoder, we employ 8 Transformer layers.
+The motion decoder consists of 12 denoising layers, including both Motion Control Layers and Temporal Align Layers.
+For the motion generator, we use 8 DiT layers.
+Throughout all modules, the hidden dimension is set to 768 and we use 8 attention heads.
+The Motion Autoencoder comprises 450 million parameters, and the motion generator comprises 200 million parameters.
+For input preprocessing, all images and video frames are resized to $256\times 256$ resolution and pixel values are normalized to the [-1,1] range.
+All models are trained under similar settings: a base learning rate of $10^{-4}$ per 64 batch size and the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\text{decay}=0.05$.
+Learning rate is set to $1e{-4}$.
+In Stage1, Models are trained with a batch size of 8 on 8 NVIDIA A800 GPUs for 300K iterations.
+In Stage2, Models are trained with a batch size of 8 on 8 NVIDIA A800 GPUs for 200K iterations.
+We employ gradient clipping with a threshold of 1.0 to stabilize training.
+Table 4: Training Configuration
+Parameter Value
+Input Video Resolution 256
+Input Video Frames 16
+Input Video FPS 25
+Optimizer Adam; $\beta_{1}=0.9,\beta_{2}=0.95$
+Learning Rate $1\times 10^{-4}$
+Warmup Steps 5000
+Learning Rate Scheduler Cosine Annealing
+$\mathcal{L}_{KL}$ 0.001
+Weight Decay 0.0001
+Training Batch Size 8
+Training Device 8 $\times$ 80G A800 GPUs
+Appendix C Userstudy
+We randomly selected 10 images from HDTF and DH-FaceVid-1K, and generated 20 videos using 2 random audio clips.
+We invited 50 participants, divided the participants into 5 groups, each participant was asked to select the best model from 4 different perspectives (realism, lip-sync, smooth, overall).
+We significantly outperform competitors in realism and match the state-of-the-art (SOTA) in lip sync, smoothness, and overall quality.
+Appendix D Additional Analysis
+D.1 Analysis on Mask
+Additional results for the cross-ID driving task are shown in fig.
+9, demonstrating that the mask strategy significantly suppresses identity leakage, although complete disentanglement is still not achieved.
+However, we emphasize that such disentanglement is not strictly necessary for video generation tasks.
+Excessive retention of fine-grained details may hinder the generation process in the second stage, while discarding unnecessary details helps stabilize both training and inference.
+As shown in  fig.
+10, even when 90% of the content is masked, the model is still able to reconstruct reasonable frames.
+This indicates a high degree of redundancy in motion information, suggesting that effective motion modeling can be achieved with only a small fraction of inputs.
+Furthermore,  fig.
+11 reveals that without the masking strategy, the model tends to memorize excessive details, which compromises its ability to generalize to out-of-distribution data.
+This observation further highlights the importance of abstract representation for robust video generation.
+D.2 More Visualization Result of Attention Map
+In fig.
+14, We visualized the attention map between the motion latent and the reference image in the decoder.
+The results show that the interaction between the motion latent and the reference image mainly focuses on facial regions, especially those areas that typically undergo significant movement.
+In addition to the face, the model also attends to local regions prone to motion, such as earrings.
+By contrast, the model pays less attention to the background and other regions that are unlikely to move.
+This phenomenon indicates that the motion latent contains rich and detailed motion information, and the model can effectively capture and utilize these highly motion-relevant local features, thereby enabling precise motion generation in target areas.
+D.3 More Visualization Result of Audio-driven Video Generation
+We provide more experimental results of talking head in fig.
+12 fig.
+13.
+In  fig.
+13, we observe that in extreme speech cases, such as silence, the labial sound ’emmm’ or certain interjections, other methods remain static, whereas our method exhibits natural behaviors, such as blinking during silence and producing natural mouth movements during interjections.
+D.4 Transfer to general video reconstruction tasks
+Since our framework is fully self-supervised, in fig.
+15, we further explored its applicability to general video scenarios.
+We found that our method is suitable for general video scenes with clear subjects and relatively simple motion.
+However, in scenarios with multiple or complex subjects, trailing artifacts may occur, as shown in the figure.
+Compared to similar methods, our approach achieves better performance on general video datasets.
+In our experiments, we reproduced the IMF method and trained it on the WebVid dataset.
+Appendix E Novelty
+The work most closely related to ours is IMF [9].
+However, our objectives and implementation strategies are fundamentally different.
+First, regarding objectives, IMF focuses solely on image-to-image transfer, whereas our goal is to transfer from images to videos.
+Moreover, the way we abstract and represent motion also differs: we employ a transformer architecture along with a masking strategy, enabling the model to encode compact and semantically rich motion features.
+Second, our decoding approach is distinct.
+We use rectified flow model as the decoder, which leads to a simpler loss design and achieves better visual results.
+Finally, we demonstrate the effectiveness of our method on the challenging downstream task of audio-driven video generation.
+Figure 9: Cross Identity Driving Example
+Figure 10: Reconstruction with Different Mask Ratio.
+Figure 11: Result of the analysis on generalization.The mask strategy enhances the model’s generalization ability.
+Figure 12: Visualization of Audio-driven Portrait Video Generation.
+Figure 13: Visualization of Audio-driven Portrait Video Generation. It can also produce natural movements—such as blinking—even during interjections or periods of silence
+Figure 14: Visualization of Attention Map.
+Figure 15: Transfer to general video reconstruction tasks.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: This paper proposes a self-supervised learning based framework for portrait motion representation, which improves the efficiency and quality of portrait video generation.
+The proposed algorithm are illustrated in Section 3.
+Experimental results are illustrated in Section 4.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We include the limitations of our work in Section 5.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: This is not a theoretical paper.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: Datasets, models, and hyperparameters used in implementing proposed algorithms are all described in detail. See Section 4 and Appendix.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We will provide the code to train and evaluate the proposed algorithm, which reproduces the experiment results in the paper after the rebuttal process.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We provide the details of our experiments in Section 4 and Appendix.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We report standard error.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We report the compute resources during training in Section 4 and Appendix.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We followed the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49. Justification: This work focuses on a academic, publicly-available benchmark. This work is not related to any private or personal data, and there’s no explicit negative social impacts.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: our work poses no risk of misuse.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: This paper cite the original papers such as dataset.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: This paper will release code for running experiments and it is well documented.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: This work does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: The LLM is used only for writing.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.